OpenStudy (anonymous):

Given cos(x)= 2/3 , with 3pi/2 < x < 2pi, find cos (x/2), sin (x/2) , and tan (x/2)

OpenStudy (jhannybean):

Well, do you know where on the unit circle \(\frac{3\pi}{2} \lt x\lt 2\pi\) is?

OpenStudy (anonymous):

that is 270 to 360

OpenStudy (jhannybean):



OpenStudy (anonymous):

got it ,Q4.

OpenStudy (jhannybean):

So, another note, do you know your half angle formulas?

OpenStudy (anonymous):

i dont know them by memory but i have them written down.

OpenStudy (jhannybean):

Can you type them here? Then I can cross check with you.

OpenStudy (anonymous):

\[\sqrt{\frac{ 1+cosA }{ 2 }} is for Cos \]

OpenStudy (anonymous):

well plus or minus in front

OpenStudy (anonymous):

should we do that one first?

OpenStudy (jhannybean):

Sure thing

OpenStudy (anonymous):

so for A would it be 2/6 plugged in?

OpenStudy (jhannybean):

I prefer writing it out in this fashion:\[\cos\left(\frac{x}{2}\right) = \pm \sqrt{\frac{1+\cos(x)}{2}}\]

OpenStudy (jhannybean):

Yes, try that out, see what you get.

OpenStudy (anonymous):

.99 and -.99

OpenStudy (anonymous):

so would it be 1 and -1?

OpenStudy (jhannybean):

Oh, I forgot to ask you one thing, is cosine positive or negative in Q4?

OpenStudy (anonymous):

oh positive

OpenStudy (anonymous):

so it has to be 1

OpenStudy (anonymous):

wait why didnt you put in "cos(2/6)" instead of "2/3"

OpenStudy (anonymous):

or is that how you check the answer? sorry, i'm horrible at this. obviously.

OpenStudy (fibonaccichick666):

ok restart

OpenStudy (fibonaccichick666):

@_lostmathstudent what does cosine mean? what is it's definition?

OpenStudy (jhannybean):

you are right, it should have been 2/6

OpenStudy (anonymous):

it's adjacent/hypotenuse

OpenStudy (fibonaccichick666):

ok so cos(2/3) what is the adj and what is the hyp?

OpenStudy (anonymous):

adj is 2 hyp is 3

OpenStudy (fibonaccichick666):

so now picture time